please answer both!!!(1 point) Use the binomial series to find the first 5 nonzero terms of the power series centered at x = 0 for the following f(1 point) For the following indefinite integral, find the full power series centered at t = 0 and then give the first 5 nonze
Answer #1

Sol 5 {1+ ] 55 I- وا 2+ 2! 8! 35 3! =i+t?)=1-49+(69=149+(294 Ginen fuction for (5+705 fo) ㅗ 55(1+2) fo (+97 0(0+1),? nontol